### 26 Mar 2026 v2

# AI generalisation gap in comorbid sleep disorder staging

Abstract
Accurate sleep staging is essential for diagnosing OSA and hypopnea in stroke patients. Although PSG is reliable, it is costly, labor-intensive, and manually scored. While deep learning enables automated EEG-based sleep staging in healthy subjects, our analysis shows poor generalization to clinical populations with disrupted sleep. Using Grad-CAM interpretations, we systematically demonstrate this limitation. We introduce iSLEEPS, a newly clinically annotated ischemic stroke dataset (to be publicly released), and evaluate a SE-ResNet plus bidirectional LSTM model for single-channel EEG sleep staging. As expected, cross-domain performance between healthy and diseased subjects is poor. Attention visualizations, supported by clinical expert feedback, show the model focuses on physiologically uninformative EEG regions in patient data. Statistical and computational analyses further confirm significant sleep architecture differences between healthy and ischemic stroke cohorts, highlighting the need for subject-aware or disease-specific models with clinical validation before deployment. A summary of the paper and the code is available at https://himalayansaswatabose.github.io/iSLEEPS_Explainability.github.io/
1 Introduction
Sleep staging is critical for diagnosing and managing sleep disorders such as obstructive sleep apnea (OSA) and hypopnea, which disrupt sleep and degrade cognition and quality of life [8]. Polysomnography (PSG), the clinical gold standard, uses EEG, EOG, and EMG to classify Wake, REM, and non-REM (N1–N3) stages, but it is costly, time-consuming, and prone to interscorer variability, especially for lighter stages like N1. Deep learning approaches provide automated and consistent alternatives with strong EEG-based performance [4, 10], yet they are trained almost exclusively on healthy data, leaving their generalization to clinical populations with disrupted sleep architecture unclear. This issue is exacerbated by the scarcity of labeled pathological data [1], highlighting the need for population-representative models and for understanding model behavior under pathology. Although sleep stage definitions should be health independent, degraded clinical performance suggests reliance on pathology-influenced or noisy features. Prior work has not modeled or interpreted sleep architecture in stroke patients with multiple comorbid sleep disorders, a group marked by severe neurological and electrophysiological abnormalities, including epilepsy, altered thalamocortical coupling, and asymmetric cortical activity, which make healthy-trained models unreliable and clinically misleading [6]. The lack of public pathology-specific EEG datasets and tailored models creates a critical gap between computational advances and clinical practice. We bridge this gap by presenting the first stroke-specific sleep staging framework, combining deep learning with medically interpretable analysis and a newly curated PSG dataset of stroke patients with comorbid sleep disorders, enabling pathology-aware and clinically deployable sleep assessment for neurologically impaired populations.
The contributions of this paper are: 1. Introduced iSLEEPS (which is accessible at https://tinyurl.com/iSLEEPSv1), a new PSG dataset of 100 ischemic stroke patients with severe sleep disorders. 2. Developed a benchmarking deep learning model for iSLEEPS that achieves state-of-the-art performance and reveals the limited generalization of models trained on healthy cohorts. 3. Statistical and explainability analyses, validated by clinician feedback, showed deep models do not generalise between healthy and comorbid cohorts, underscoring the need for subject-aware or disease-specific sleep staging models.
2 Methodology
2.1 Model Architecture
The model (Fig. 1) uses a sliding window of consecutive 30 s EEG epochs to capture temporal context and predicts the sleep stage of the central epoch. A SE-ResNet block extracts discriminative spectral–temporal features by enhancing relevant frequency–amplitude patterns while suppressing noise, while stacked Bi-LSTM layers model bidirectional temporal dependencies and long-range stage transitions. The encoded features are then passed through fully connected layers, with the final layer producing sleep stage class probabilities.
2.2 Explainability Methods
We used GradCAM [9] for explainability. Gradient-weighted Class Activation Mapping (Grad-CAM) generates class-specific heatmaps from the gradients of a target class with respect to the final convolutional layer. In signal classification, it highlights time or frequency regions most influential to the model’s decisions, enabling assessment of whether the model attends to meaningful physiological features (e.g., sleep spindles or alpha rhythms), thereby improving interpretability, supporting debugging, and ensuring alignment with domain knowledge.
3 Experiments
We evaluated performance using raw single-channel EEG (C4:M1) from SleepEDF-20 [3], SleepEDF-78, SHHS [12], and our iSLEEPS dataset. SleepEDF-20 includes 39 PSG recordings from 20 healthy adults aged 25–34 with R&K staging (W, N1, N2, N3 [N3+N4], REM), excluding Movement and UNKNOWN stages. SleepEDF-78 extends this to 78 participants. SHHS provides full-night PSG data from 6,441 subjects scored under AASM guidelines, with analysis restricted to 329 typical sleepers. iSLEEPS comprises 100 ischemic stroke patients (mean age 50.52; 23 females, 77 males) recruited at NIMHANS, Bangalore, India. It shows a high rate of sleep apnea (38% severe, 23% moderate) with detailed obstructive/central apnea annotations Deep learning models were trained separately per dataset using Adam optimizer(learning rate 0.001), negative log-likelihood loss, batch size 128, and evaluated via leave-one-out cross-validation (20, 10, 5, and 10 folds for SleepEDF-20, SleepEDF-78, SHHS, and iSLEEPS, respectively). The architecture processes raw EEG with a window size 9, stride 4, and three stacked LSTM layers, with the best healthy-subject model tested on iSLEEPS to simulate clinical deployment, followed by GradCAM analysis on five randomly selected patients to assess errors and alignment with clinically relevant EEG patterns.
4 Results and Discussion
4.1 Model Performance
| Datasets | Model | Overall Results | ||
| ACC | MF1 | |||
| SleepEDF-20 (Healthy Cohort) | NAS [4] | 82.7 | 75.9 | 0.76 |
| DeepSleepNet [10] | 82.0 | 76.9 | 0.76 | |
| XSleepNet2 [7] | 83.9 | 78.7 | 0.77 | |
| SleepContextNet [13] | 84.8 | 79.8 | 0.79 | |
| SEResnet-Transformer [5] | 79.3 | 74.7 | 0.72 | |
| Ours | 87.5 | 82.5 | 0.82 | |
| SleepEDF-78 (Healthy Cohort) | NAS [4] | 80.0 | 72.7 | 0.72 |
| XSleepNet2 [7] | 80.3 | 76.4 | 0.73 | |
| SleepContextNet [13] | 82.7 | 77.2 | 0.76 | |
| SEResnet-Transformer [5] | 73.6 | 70.6 | 0.68 | |
| Ours | 83.8 | 78.9 | 0.77 | |
| SHHS (Healthy Cohort) | NAS [4] | 81.9 | 75.3 | 0.74 |
| SleepContextNet [13] | 86.4 | 80.5 | 0.81 | |
| SEResnet-Transformer [5] | 79.0 | 69.3 | 0.71 | |
| Ours | 87.8 | 81.9 | 0.83 | |
| iSLEEPS (Patient Cohort) | ResNet [2] | 61.6 | 54.4 | 0.48 |
| SEResnet-Transformer [5] | 67.4 | 59.3 | 0.54 | |
| Ours | 74.7 | 67.7 | 0.64 | |
We provide a comprehensive comparison with the current literature in Table 1 and an overview of the model’s sleep stage-wise performance across four different datasets in Table 2. This validates the fact the proposed benchmarking model on iSLEEPS is at par and, in fact, better than the existing models in the current literature for sleep staging in the available healthy subject datasets, as well as in patient-based cohorts.
| Dataset | W | N1 | N2 | N3 | REM |
| SleepEDF-20 | 92.0 | 56.9 | 89.9 | 87.9 | 85.9 |
| SleepEDF-78 | 92.1 | 50.1 | 85.2 | 82.1 | 82.1 |
| SHHS | 89.1 | 54.4 | 89.2 | 87.9 | 88.9 |
| iSLEEPS | 79.9 | 32.9 | 80.9 | 74.2 | 70.6 |
4.2 Ablation studies
We performed ablation studies on the SleepEDF-20 and iSLEEPS datasets using 20-fold and 10-fold cross-validation, respectively, to assess the impact of key model components. A window size of 9 with a stride of 4 yielded the best balance between temporal coverage and efficiency. SE-ResNet-18 outperformed SE-ResNet-34 as the feature extractor, with the latter showing overfitting, and the SE module significantly improved feature learning. Optimal performance was achieved with SE-ResNet-18 and three Bi-LSTM stacks, as additional stacks increased complexity without improving results.
4.3 Model Explainability and Clinician Feedback
For the 100 iSLEEPS patients, sleep stages were evaluated using the SHHS-trained model, yielding 55.1% accuracy and 51.9% MF1, a sharp drop from Table 1. To interpret its behavior, GradCAM heatmaps were generated for all epochs of four randomly selected patients. Medical experts found that, despite training on healthy data, the model often attended to physiologically irrelevant EEG regions in stroke patients. Although some epochs showed clinically meaningful patterns, many activations reflected diffuse slowing, intermittent delta bursts, or movement and electrode artifacts. They did not match typical sleep markers such as spindles, K-complexes, or alpha attenuation. This mismatch was most pronounced during N1–N2 transitions and REM detection, where attention favored ischemia-related slowing or hemispheric asymmetries over lesioned areas.
Overall, the model generalizes poorly from healthy subjects to patients, frequently misdirecting attention, missing clinically relevant cues, and producing systematic staging errors. Figure 2 examines two representative epochs (trained on healthy data, tested on patients) using medical annotations and GradCAM. In Figure 2(a), the prediction is correct but attention is concentrated on physiologically irrelevant regions, indicating poor medical alignment. In Figure 2(b), attention is closer to relevant regions, yet the prediction is incorrect, again highlighting a disconnect between clinical reasoning and model behavior. By contrast, a model trained and evaluated on iSLEEPS (Table 1) shows improved focus on clinically meaningful features (Figure 3), suggesting superior capture of patient-specific sleep architecture.
These results indicate that models trained on healthy-subject datasets (SleepEDF-20, SleepEDF-78, SHHS) should be applied in clinical or diagnostic settings only under medical supervision. This limitation is expected, as sleep architecture differs substantially between ischemic stroke patients and healthy individuals [6], a disparity our experiments explicitly expose through the observed performance degradation.
4.4 Clincian-in-the-loop Analysis of Sleep Transition Graphs
We isolated epochs from SleepEDF-78 and iSLEEPS to form a balanced dataset containing 22,430 (49%) epochs from SleepEDF-78 and 23,192 (51%) from iSLEEPS, ensuring substantial representation of both healthy and patient samples. Significance testing confirmed that the sleep architectures of the two groups differ notably in stage frequency and order. The state-transition diagrams (Fig 4) also show significant differences (diagram excludes insignificant transitions and self-transitions), with an average KL Divergence of , a Chi-Square Statistic of with 19 degrees of freedom, and p-values ranging from to across 13 significant transitions out of 20 (Mean p-value , Median p-value ). The analysis in this section was performed in direct collaboration with our clinician co-author.
The Statistical Lens: Transition probability analysis revealed clear differences between stroke patients (iSLEEPS) and healthy individuals (SleepEDF-78). Stroke patients showed a higher likelihood of waking from light non-REM sleep (N2Wake: 0.354 vs. 0.164), indicating unstable sleep marked by frequent arousal rather than progression into deep restorative sleep. Clinically, this instability reduces sleep efficiency and restorative benefits, consistent with post-stroke impairment of thalamocortical and brainstem systems regulating sleep and arousal. Stroke patients showed slower N2N3 transitions, suggesting disrupted sleep depth. Healthy individuals showed stronger transitions, consistent with stable and consolidated sleep. These findings align with prior studies showing that cortical and subcortical brain lesions disrupt neural networks and lead to sleep disturbances after stroke [6].
The Computational Lens: Given clear differences between patient and healthy samples, we trained binary classifiers (Random Forest, Logistic Regression, Decision Tree, SVM) to predict whether an unepoched EEG recording came from a patient or a healthy subject. Using targeted feature engineering, we extracted 27 structural and sequential features; One of such features which perfectly separated the cohorts was Average Run Length (a Run is defined as a continuous segment without stage change) as seen in Figure 5. Combinations of bigrams (e.g., WakeN1 with N3N2), Shannon entropy of the arrangement of the sleep stages, and maximum Run length also achieved 100% accuracy with zero AUC-ROC variance across five folds. These findings show that ischemic stroke patients have markedly distinct sleep architecture, explaining poor generalization from healthy-only models and highlighting the need for domain-aware clinical sleep staging.
5 Conclusion
Models trained only on healthy sleep data perform poorly on patient sleep staging, showing an 30% accuracy drop and strong dataset bias, and no public benchmarks exist for multimorbid sleep disorders. We highlight two urgent needs: models that handle inter-patient and structural variability, such as hierarchical approaches separating healthy and patient data, and mandatory expert review before clinical use due to reliance on non-pathological features. To address this gap, we introduce iSLEEPS, a forthcoming public benchmark, noting that medical-grade, explainable, and unbiased sleep staging remains an open challenge under high-risk Health AI and the EU AI Act [11].